Brief Exercise 10-10o For its three investment centers, Gerrard Company accumulates the following data: Sales Controllable margin Average operating assets the centers expect the folio ing changes n the next year: 1 increase sales 20%, a decrease costsp90,000: Compute the expected return on investment (ROI) for each center. Assume center t has a controlable margin percentage of 70% (Round nor to 1 deci $1,980,000 $3,916,000 $4,041,000 1,169,320 2,150,280 4,597,620 5,084,000 7,964,000 12.099,000 (III) decrease average operating assets $496,000. al p ce, ag. I SJ
Answer #1
Calculation of return on investment:
Return on investment= controllable margin/average operating assets*100
I)
Sales=$1980000
Contribution margin= 1980000*0.70=$1386000
Controllable Fixed cost= Contribution margin- controllable margin
                                              =1386000-1169320=$216680
Sales (1980000*1.20)=$2376000
Contribution margin= 2376000*0.70=$1663200
Controllable margin=1663200-216680=$1446520
Average operating assets=$5084000
Return on investment= 1446520/5084000*100=28.5%
Expected Return on investment=28.5%
II)
Cotrollable margin= 2150280+390000=$2540280
Average operating assets= $7964000
Return on investment= 2540280/7964000*100=31.9%
Exoected Return on investment= 31.9%
III)
Controllable margin=$4597620
Average operating assets=12099000-496000=$11603000
Return on investment= 4597620/11603000*100=39.6%
Expected Return on Investment=39.6%